OpenStudy (anonymous):

integrate cos^2(2x) sin(2x)

OpenStudy (helder_edwin):

u mean \[ \large \int\cos^2(2x)\ \sin(2x)\,dx \]?

OpenStudy (helder_edwin):

use the substitution \[ \large u=\cos(2x) \]

OpenStudy (anonymous):

I get -(1/6)cos^3(2x)

OpenStudy (helder_edwin):

let's see \[ \large u=\cos(2x)\qquad\text{then}\quad du=-2\sin(2x) \]

OpenStudy (helder_edwin):

so \[ \large \int\cos^2(2x)\sin(2x)\,dx=\int u^2\frac{du}{-2} \] \[ \large =\frac{1}{-2}\frac{u^3}{3} \] so \[ \large \int\cos^2(2x)\sin(2x)\,dx=-\frac{1}{6}\cos^3(2x)+C \]